OpenStudy (anonymous):

Suppose a classmate claims that the difference between (x + 2)2 and (x2 + 4) must be 0. Is your classmate correct? Choose the best explanation. A. Yes; (x + 2)2 = x2 + 4; x2 + 4 – (x2 + 4) = 0. B. Yes; when you substitute –2 for x, (–2 + 2)2 = 0 and (–2)2 + 4 = 0. C. No; (x + 2)2 = (x2 + 4x+ 4) – (x2 + 4) = 4x + 8. D. No; (x + 2)2 = x2 + 4x + 4 look at the picturex2 + 4.

OpenStudy (anonymous):

I think the answer has to be no, but why

OpenStudy (anonymous):

the part that says "look at the picture" is supposed to say "not equal"

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

is it (x+2)^2 ???

OpenStudy (anonymous):

is what (x+2)^2?

OpenStudy (anonymous):

thats not true though, they dont equal eachother

OpenStudy (anonymous):

\[x^{2} + 4x + 4 = (x+2)^{2}\]

OpenStudy (anonymous):

(x+2)^2 = x^2 +4x+4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so, ya, its not true. x^2 +4x+4 is not equel to x^2 +4 unless x=0

OpenStudy (anonymous):

but, 2 answers say its not true, it is c or d

OpenStudy (anonymous):

its c. because thesubstracted result is not equel to zero, that one agrees with question

OpenStudy (anonymous):

lol it was wrong, the answer was c... oh well